OpenStudy (unklerhaukus):

\[ \mathcal L ^{-1}\left\{\frac{e^{-\pi p}}{p^2+1}\right\}\]

OpenStudy (unklerhaukus):

\[ \begin{align*} \mathcal L ^{-1}\left\{\frac{e^{-\pi p}}{p^2+1}\right\} &=h(t-\pi)\times\mathcal L^{-1} \left\{\frac1{p^2+1}\right\}_{t\rightarrow t-\pi}\\ &=h(t-\pi)\times\sin(t)\Big|_{t\rightarrow t-\pi}\\ &=h(t-\pi)\times\sin(t-\pi)\\ &=-h(t-\pi)\sin(t)\\ \end{align*}\]

OpenStudy (unklerhaukus):

@AccessDenied

OpenStudy (unklerhaukus):

have i got this one right/?

OpenStudy (accessdenied):

This appears correct to me. :)

OpenStudy (unklerhaukus):

thanks!, would you more likely leave the answer in the final form of \[=h(t−π)\sin(t−π)\]or \[=−h(t−π)\sin(t)\]

OpenStudy (accessdenied):

I would probably use -h(t-pi) sin(t) just because I prefer to see the simplest forms of functions..